A 300f capacitor separated by a distance of 4.0mm was charged to a potential difference of 250 V. What is the energy density in its region in terms of jouled/m³


1
Expert's answer
2022-03-28T11:02:51-0400

The electric field between plates

"E=\frac{V}{d}=\frac{250}{0.004}=6.25*10^4\:\rm V/m"

The energy density

"w=\frac{\epsilon_0E^2}{2}\\
=\frac{8.85*10^{-12}*(6.25*10^4)^2}{2}=0.017\:\rm J"